OpenStudy (anonymous):

indices= simplify 9^1 and a 1/2 pls help

OpenStudy (anonymous):

@Linchpin @confluxepic @TheSmartOne

OpenStudy (confluxepic):

They are simplified.

OpenStudy (confluxepic):

You can't further simplify 1/2.

OpenStudy (confluxepic):

Oh. Wait. Indices.

OpenStudy (campbell_st):

rewrite the power as an improper fraction \[9^{\frac{3}{2}}\] and rewrite 9 as a power of 3 then use the power of a power law... multiply the powers \[(3^2)^{\frac{3}{2}} = 3^{\frac{3}{2} \times 2}\] and simplify

OpenStudy (anonymous):

does it equal to 9 over 2??

OpenStudy (anonymous):

is the final answer 9 --- 2 @campbell_st

OpenStudy (campbell_st):

I'm not sure... as I've answered this question simplify \[9^{~1\frac{1}{2}}\] if thats the question you need to simplify \[3^{\frac{3}{2} \times 2}\]

OpenStudy (anonymous):

@campbell_st would the answer be 27?

OpenStudy (campbell_st):

that's correct